# Learning image representations for anomaly detection: application to discovery of histological alterations in drug development

### 02 Dec 2023 v6

@@ 1 Introduction
Figure 1: Anomaly detection approach. A: Learning image representations with an auxiliary supervised classification task on set of tissue categories. The figure on the right shows t-SNE plot of image features after the training. Clusters marked with arrows correspond to target classes for the anomaly detection task in our experiments. Each color corresponds to a particular category (a combination of specie, organ, and staining). B: Training one-class classifier on a tissue of a category of interest. C: Anomaly detection in tissue of the chosen (target) category. Anomaly score $\alpha$ can be thesholded to output a binary decision.
@@ 3.1.1 CNN encoder: learning image representations
The critical part of the AD system is a CNN encoder that should generate feature vectors with high representation power allowing discrimination between anomalous and normal tissue structures. On the other, representation vectors should be short enough to generalize to unseen data. Preferably, to facilitate operation of a one-class classifier, representations should be invariant to variations in image appearance due to differences in staining procedure and image acquisition. CNNs that were pre-trained on ImageNet dataset may generate feature vectors that do not have the mentioned above properties. We therefore propose a few techniques for training CNN to adapt image representations to histopathlogical images.
@@ 3.1.4 Objective function
Training the CNN on the auxiliary task with a cross-entropy loss objective function will force learning image representations that are discriminative for histological structures inherent to the species, organs, and stainings present in the used dataset. We cannot, however, train them to be disciminative for anomalies, since those are not available or even not yet known. One of the approaches for AD is to minimize the volume of hypersphere in the feature space where majority of normal data lives (Tax and Duin, 2004; Ruff et al., 2018). Conceptually similar, in Perera and Patel (2019) compactness loss was used to keep low intra-class variance in feature space for normal data. We will also enforce image representations to be compact, while training CNN image representations on the auxiliary task (Fig. 1 Top). The compact representations should increase the sensitivity of the system to anomalies. For this purpose we use center-loss $\mathcal{L}_{CL}$ term (Wen et al., 2016) in the objective function, along with the multiclass cross-entropy $\mathcal{L}_{CE}$ term
@@ 4 Experiments
We implemented all the experiments in the a PyTorch deep learning framework (Paszke et al., 2019). Trained models and code are available online [^5] Boehringer-Ingelheim-anomaly-detection-in-histology.
@@ 4.2 Training image representations
To train image representations using an auxiliary task as described in Sec. 3.1.2 (See Top of Fig. 1) we collected a large dataset of histological $256\times 256$, 0.442 µm/pix. tiles extracted from Whole Slide Images (WSI) of healthy animals tissue. Seven organs (liver, brain, kidney, heart, lung, pancreas, spleen), two animal species (rat and mouse), and two types of staining (Masson Trichrome and H&E) comprises the collected dataset. Overall we have 16 categories. Note that not all combinations of specie, organ, and staining was available for us (rat tissue was available only for liver). Around 7000 tiles from each of the 16 categories is sampled from the dataset. During the training of FC-NN and CNN models and we will use 6 different random seeds that randomizes the choice of the 7000 sampled tiles and the order of their feeding to the network. During the test time we will provide mean and standard error values over 6 trained models, which will allow us more reliable performance comparisons. Note that the extracted from WSI tiles were not curated by a human and may occasionally contain various artifacts or wrong tissue. We made this dataset of healthy tissue publically available (Zingman et al., 2022) such that training representations of histological images for our AD approach can be reproduced.
@@ 4.5 Performance evaluation based on NAFLD anomalies
We evaluate the performance of our AD system (see, Fig. 1), C) using encoders based on a few different backbone CNN architectures. Here we report the performance using the balanced accuracy and provide the mean over the results obtained from six trained CNN encoders that were trained with six seeds (0, 100, 200, 300, 400, 500). We also provide a standard error for the estimated means. The results are summarized in Table 1 for both Masson Trichrome and H&E staining. The average between these performances is reported in the last row.
@@ 4.5.1 Advantage of learning representations for histological images
Here we provide a comparison of our AD system, which is composed of a CNN encoder and a one-class classifier, with the system of a similar architecture but with the CNN encoder trained on ImageNet without further adaptation to histopatholgical data using our auxiliary task. Table 2 summarizes the performances of AD for a set of CNN architectures (the same set as in Sec.  4.5) trained on the ImageNet dataset. Put attention that in this table we report balanced accuracy measures with no standard error as in Sec. 4.5, where we also reported the standard error of the estimated mean over a few CNN encoders trained on random subsets of histological images.
Comparing Tables 1 and 2 one can see an essential increase in the performance when CNN encoders were trained on histological data. For example, for the best performing architecture EfficientNet-B0-320 the balanced accuracy increases in $26\%$. The results emphasize the importance of tuning the neural networks to a particular target domain. Even though recently it was shown that using transfer learning concept for AD (Bergmann et al., 2020; Rippel et al., 2020), where ImageNet trained neural networks encode images with feature vectors representations, can achieve state of the art results. However, for many cases where target domain diverges from natural images, like in our case of histopatohlogical images, tuning neural networks to the target domain is critical.
@@ 4.5.3 Comparison with self supervised learning approaches
We also compared feature representations obtained with SSL, ImageNet pre-trained neural network, and supervised learning on our auxiliary task, when all the methods used the neural network of the same ResNet-18 architecture and the anomaly detection performance was measured using balanced accuracy [^14] For SimCLR and SimSiam we set margin error $\nu$ of one-class SVM to 0.3, which gave us the best performance for the SSL methods.. From Table 5 it can be seen that SSL was able to improve performance over the ImageNet pre-trained network, but falls far behind the performance of the features learned from our auxiliary task. The results show that with readily available labels for histopathogical data from different organs, species, and stained with different reagents, fully supervised training allows to learn powerful features that are considerably better then when using self-supervised learning techniques. Both contrastive and non-contrastive self-supervision forces feature similarity for images after applied transformations, but not similarity and dissimilarity for images within and between histopathological tissue classes.
@@ 4.8 Conclusion
We developed an approach enabling detection of abnormal structures in histopathological images under the real-world environment. We introduced a recipe for training image representations adapting them to the domain of histopathology, which allowed a standard anomaly detection to achieve strong performance. We collected and published a dataset of histopathologcal images of tissue with non-alcoholic fatty liver disease (NAFLD) that exhibit different abnormal variations and can be used as a benchmark for anomaly detection in histological images. Using this dataset we have shown that our approach outperforms the recent state-of-the-art and classical methods for anomaly detection. Moreover, we have shown that our method behaves similarly to established methods designed specifically to quantify NAFLD condition. We have also demonstrated that our approach was able to reveal tissue alteration in the liver resulted from side effects of a drug under development. We can conclude that the developed approach is capable of detection of unknown adverse drug effects and has a potential of reducing attrition rates in the drug development.
@@
Currently, our system cannot use anomaly examples to improve its performance. Moreover, re-training one-class classifier on continuously gathered new normal data won’t be efficient until the amount of the new data is relatively large. In future, we, therefore, aim at extending our anomaly detection approach to be able to gradually improve performance using both anomalous and normal data that occasionally becomes available, i.e moving to continual learning regime.

### 08 Jan 2024 v7

@@ 1 Introduction
Figure 1: Anomaly detection approach. A: Learning image representations with an auxiliary supervised classification task on set of tissue categories. The figure on the right shows t-SNE plot of image features after the training. Clusters marked with arrows correspond to target classes for the anomaly detection task in our experiments. Each color corresponds to a particular category (a combination of specie, organ, and staining). B: Training one-class classifier on a tissue of a category of interest. C: Anomaly detection in tissue of the chosen (target) category. Anomaly score $\alpha$ can be thresholded to output a binary decision.
@@ 3.1.1 CNN encoder: learning image representations
The critical part of the AD system is a CNN encoder that should generate feature vectors with high representation power allowing discrimination between anomalous and normal tissue structures. On the other, representation vectors should be short enough to generalize to unseen data. Preferably, to facilitate operation of a one-class classifier, representations should be invariant to variations in image appearance due to differences in staining procedure and image acquisition. CNNs that were pre-trained on ImageNet dataset may generate feature vectors that do not have the mentioned above properties. We therefore propose a few techniques for training CNN to adapt image representations to histopathological images.
@@ 3.1.4 Objective function
Training the CNN on the auxiliary task with a cross-entropy loss objective function will force learning image representations that are discriminative for histological structures inherent to the species, organs, and stainings present in the used dataset. We cannot, however, train them to be discriminative for anomalies, since those are not available or even not yet known. One of the approaches for AD is to minimize the volume of hypersphere in the feature space where majority of normal data lives (Tax and Duin, 2004; Ruff et al., 2018). Conceptually similar, in Perera and Patel (2019) compactness loss was used to keep low intra-class variance in feature space for normal data. We will also enforce image representations to be compact, while training CNN image representations on the auxiliary task (Fig. 1 Top). The compact representations should increase the sensitivity of the system to anomalies. For this purpose we use center-loss $\mathcal{L}_{CL}$ term (Wen et al., 2016) in the objective function, along with the multiclass cross-entropy $\mathcal{L}_{CE}$ term
@@ 4 Experiments
We implemented all the experiments in the PyTorch deep learning framework (Paszke et al., 2019). Trained models and code are available online [^5] Boehringer-Ingelheim-anomaly-detection-in-histology.
@@ 4.2 Training image representations
To train image representations using an auxiliary task as described in Sec. 3.1.2 (See Top of Fig. 1) we collected a large dataset of histological $256\times 256$, 0.442 µm/pix. tiles extracted from Whole Slide Images (WSI) of healthy animals tissue. Seven organs (liver, brain, kidney, heart, lung, pancreas, spleen), two animal species (rat and mouse), and two types of staining (Masson Trichrome and H&E) comprises the collected dataset. Overall we have 16 categories. Note that not all combinations of specie, organ, and staining was available for us (rat tissue was available only for liver). Around 7000 tiles from each of the 16 categories is sampled from the dataset. During the training of FC-NN and CNN models and we will use 6 different random seeds that randomizes the choice of the 7000 sampled tiles and the order of their feeding to the network. During the test time we will provide mean and standard error values over 6 trained models, which will allow us more reliable performance comparisons. Note that the extracted from WSI tiles were not curated by a human and may occasionally contain various artifacts or wrong tissue. We made this dataset of healthy tissue publicly available (Zingman et al., 2022) such that training representations of histological images for our AD approach can be reproduced.
@@ 4.5 Performance evaluation based on NAFLD anomalies
We evaluate the performance of our AD system (see Fig. 1C) using encoders based on a few different backbone CNN architectures. Here we report the performance using the balanced accuracy and provide the mean over the results obtained from six trained CNN encoders that were trained with six seeds (0, 100, 200, 300, 400, 500). We also provide a standard error for the estimated means. The results are summarized in Table 1 for both Masson Trichrome and H&E staining. The average between these performances is reported in the last row.
@@ 4.5.1 Advantage of learning representations for histological images
Here we provide a comparison of our AD system, which is composed of a CNN encoder and a one-class classifier, with the system of a similar architecture but with the CNN encoder trained on ImageNet without further adaptation to histopathological data using our auxiliary task. Table 2 summarizes the performances of AD for a set of CNN architectures (the same set as in Sec.  4.5) trained on the ImageNet dataset. Put attention that in this table we report balanced accuracy measures with no standard error as in Sec. 4.5, where we also reported the standard error of the estimated mean over a few CNN encoders trained on random subsets of histological images.
Comparing Tables 1 and 2 one can see an essential increase in the performance when CNN encoders were trained on histological data. For example, for the best performing architecture EfficientNet-B0-320 the balanced accuracy increases in $26\%$. The results emphasize the importance of tuning the neural networks to a particular target domain. Even though recently it was shown that using transfer learning concept for AD (Bergmann et al., 2020; Rippel et al., 2020), where ImageNet trained neural networks encode images with feature vectors representations, can achieve state of the art results. However, for many cases where target domain diverges from natural images, like in our case of histopathological images, tuning neural networks to the target domain is critical.
@@ 4.5.3 Comparison with self supervised learning approaches
We also compared feature representations obtained with SSL, ImageNet pre-trained neural network, and supervised learning on our auxiliary task, when all the methods used the neural network of the same ResNet-18 architecture and the anomaly detection performance was measured using balanced accuracy [^14] For SimCLR and SimSiam we set margin error $\nu$ of one-class SVM to 0.3, which gave us the best performance for the SSL methods.. From Table 5 it can be seen that SSL was able to improve performance over the ImageNet pre-trained network, but falls far behind the performance of the features learned from our auxiliary task. The results show that with readily available labels for histopathological data from different organs, species, and stained with different reagents, fully supervised training allows to learn powerful features that are considerably better then when using self-supervised learning techniques. Both contrastive and non-contrastive self-supervision forces feature similarity for images after applied transformations, but not similarity and dissimilarity for images within and between histopathological tissue classes.
@@ 4.8 Conclusion
We developed an approach enabling detection of abnormal structures in histopathological images under the real-world environment. We introduced a recipe for training image representations adapting them to the domain of histopathology, which allowed a standard anomaly detection to achieve strong performance. We collected and published a dataset of histopathological images of tissue with non-alcoholic fatty liver disease (NAFLD) that exhibit different abnormal variations and can be used as a benchmark for anomaly detection in histological images. Using this dataset we have shown that our approach outperforms the recent state-of-the-art and classical methods for anomaly detection. Moreover, we have shown that our method behaves similarly to established methods designed specifically to quantify NAFLD condition. We have also demonstrated that our approach was able to reveal tissue alteration in the liver resulted from side effects of a drug under development. We can conclude that the developed approach is capable of detection of unknown adverse drug effects and has a potential of reducing attrition rates in the drug development.
@@
Currently, our system cannot use anomaly examples to improve its performance. Moreover, re-training one-class classifier on continuously gathered new normal data will not be efficient until the amount of the new data is relatively large. In future, we, therefore, aim at extending our anomaly detection approach to be able to gradually improve performance using both anomalous and normal data that occasionally becomes available, i.e moving to continual learning regime.